OpenStudy (anonymous):

utilizando 5 como el valor de x, encontrar y en la ecuacion: -3x + 2y = 1

OpenStudy (mathstudent55):

Haga lo que dice, utilize el valor the 5 en x: -3(5) + 2y = 1 -15 + 2y = 1 +15 +15 ----------------- 2y = 16 y = 8

OpenStudy (anonymous):

graciaz

OpenStudy (mathstudent55):

de nada